Evaluate integral integral integral yz dv where E is the region bounded by x = 2y^2 + 2z^2 - 5 and the plane x = 1
1
Expert's answer
2021-07-22T16:48:36-0400

Answer:-

We’ll integrate in the order "dxdydz."



"x=1: 1=2y^2+2z^2-5""y^2+z^2=3"

We have "Q=\{(y, z):y^2+z^2\leq3\}" and



"\int \int \int_E yzdV=\int \int_Q\bigg(\displaystyle\int_{2y^2+2z^2-5}^{1}yzdx\bigg)dA""=\int \int_Q(1-2y^2-2z^2+5)yzdA""=\displaystyle\int_{0}^{2\pi}\displaystyle\int_{0}^{\sqrt{3}}(6-2r^2)r\sin\theta r\cos\theta rdrd\theta""=\displaystyle\int_{0}^{2\pi}\sin (2\theta )\bigg[\dfrac{3r^4}{4}-\dfrac{r^6}{6}\bigg]\begin{matrix}
 \sqrt{3} \\
 0
\end{matrix}d\theta""=\displaystyle\int_{0}^{2\pi}\dfrac{9}{4}\sin (2\theta )d\theta""=-\dfrac{9}{8}[\cos(2\theta)]\begin{matrix}
 2\pi \\
 0
\end{matrix}=0"